# A Note on the Legendre Transformation

### v1

@@ 1 Curves and singularities
It is proved that for $n=2$ or $3$ the germ of any curve (3) at zero can be brought to the simplest form $x=t^{n}$, $y=t^{n+1}$ by means of an appropriate change of variables $(x,y)$ and a change of the parameter $t$. However, for $n\geq 4$ it is not true [5].
@@ 2.1 Legendrian lift
The lift or $1$-graph of the curve $\gamma$ given by formula (1) is the curve
@@
The $1$-graph of the curve $\gamma$ does not depend on the choice of the parametrization the curve $\gamma$.
According to modern biological concepts going back to the famous work of Hubel and Wiesel (1959), the brain of mammals performs the Legendrian lift of apparent contours of flat images perceived by the retina. There exists an area of the primary cortex of the brain associated with vision (V1) that contains neurons united in groups with complex hierarchy (columns, hypercolumn). Each column is sensitive to a certain point of the retina and a certain tangent direction at this point. Flat images perceived by the retina are lifted in the space of 1-jets, and further work of the brain occurs with the lifted objects. See [12] and the references therein. The only difference between the lift performed by the brain and the lift described above is that the brain uses the third coordinate $\theta=\arctan p$ instead of $p$. It is very natural: $\theta$ is more appropriate for procedure performed biologically, since it is bounded, while $p$ is not.
@@ 2. Assume that the curve $\gamma$ is the graph of a functions $y=f(x)$ such that
and if $f$ is strictly concave, its Legendre transformation $f^{*}$ can be defined by the formula
@@
where $R$ denoted the range of $f^{\prime}$.
@@
The Legendre transformation can be obviously adapted for hypersurfaces and, consequently, for functions depending on any number of variables. In particular, this way brings us to the same formulas (8) and (9), where $x$ and $p$ belong to the spaces of the same dimension and $xp$ means the dot product in orthonormal coordinates.
@@ 3 Contact transformations
To define pedal transformation, let us fixe an arbitrary point $O$ on the plane (without loss of generality, $O$ can be the origin). The point $O$ is called the pole or center of the transformation. Let $\gamma$ be a curve on the plane. Then the pedal curve of $\gamma$ is the locus of points $X$ so that the line $OX$ is perpendicular to the tangent line to $\gamma$ passing through the point $X$.

### v2

@@ 1 Curves and singularities
It is proved that for $n=2$ or 33 the germ of any curve (3) at zero can be brought to the simplest form $x=t^{n}$, $y=t^{n+1}$ by means of an appropriate change of variables $(x,y)$ and a change of the parameter $t$. However, for $n\geq 4$ it is not true [5].
@@ 2.1 Legendrian lift
The lift or 11-graph of the curve $\gamma$ given by formula (1) is the curve
@@
The 11-graph of the curve $\gamma$ does not depend on the choice of the parametrization the curve $\gamma$.
According to modern biological concepts going back to the famous work of Hubel and Wiesel (1959), the brain of mammals performs the Legendrian lift of apparent contours of flat images perceived by the retina. There exists an area of the primary cortex of the brain associated with vision (V1) that contains neurons united in groups with complex hierarchy (columns, hypercolumn). Each column is sensitive to a certain point of the retina and a certain tangent direction at this point. Flat images perceived by the retina are lifted in the space of 1-jets, and further work of the brain occurs with the lifted objects. See [12] and the references therein. The only difference between the lift performed by the brain and the lift described above is that the brain uses the third coordinate $\theta=\arctan p$ instead of pp. It is very natural: $\theta$ is more appropriate for procedure performed biologically, since it is bounded, while $p$ is not.
@@ 2. Assume that the curve $\gamma$ is the graph of a functions $y=f(x)$ such that
see Fig. 5. If the function $f$ is strictly concave, its Legendre transformation $f^{*}$ can be defined by the formula
@@
where in the both formulas $R$ denotes the range of $f^{\prime}$.
Figure 5: Illustration of formula (8). On the right: the case that $p\notin R$.
Let us prove formula (8) (for (9) the proof is similar). Fixing arbitrary $p\in R$, consider the function F(x):=xpf(x)F(x):=xp-f(x). Since the function $f$ is strictly convex, $F$ is strictly convex as well. Therefore, $F$ has no more than one critical point.
Assume that $F$ has a critical point, denote it by $t$ (Fig. 5 left). Then $F^{\prime}(t)=0$ and $f^{\prime}(t)=p$, whence from (8) we get:
$f^{*}(p)=\sup\limits_{x}\,(xp-f(x))=tp-f(t)=tf^{\prime}(t)-f(t),\ \ p=f^{\prime}(t).$ (10)
The graph of the function $f^{*}(p)$ given by formula (10) coincides with the Legendre transformation of the graph $y=f(x)$. Indeed, let us introduce the parameter $t=x$, then $y=f(t)$, $p=f^{\prime}(t)$, This yields
$X=f^{\prime}(t),\ \ \,Y=tf^{\prime}(t)-f(t),$
that is, (10) is equivalent to $f^{*}(X)=Y$.
If for the chosen $p$ the function $F$ has no critical point (Fig. 5 right), the value $f^{*}(p)$ cannot be defined by either the first or second way. However, in such case one can set $f^{*}(p)=+\infty$ according to (8).
@@
The Legendre transformation can be obviously adapted for hypersurfaces and, consequently, for functions depending on any number of variables. In particular, this way brings us to the same formulas (8) and (9), where $x$ and $p$ belong to the spaces of the same dimension and xpxp means the dot product in orthonormal coordinates.
@@ 3 Contact transformations
To define pedal transformation, let us fixe an arbitrary point $O$ on the plane (without loss of generality, $O$ can be the origin). The point $O$ is called the pole or center of the transformation. Let $\gamma$ be a curve on the plane. Then the pedal curve of $\gamma$ is the locus of points $X$ so that the line OXOX is perpendicular to the tangent line to $\gamma$ passing through the point $X$.